# How Does a Deep Neural Network Look at Lexical Stress?

### v2

@@ I Introduction
While deep learning based architectures have transformed many aspects of speech processing, those models often exhibit an nontransparent prediction process, obscuring how specific decisions are made. This work explores the application of one technique for shedding light on such architectures through the lens of a well-studied phenomenon in speech science and speech processing: lexical stress. Lexical stress refers to the prominence, accentuation, or emphasis of a particular syllable within a word. In many languages, lexical stress is crucial for communication as it can serve as the primary basis for a contrast in meaning. For example, in English, the words “PROtest” (noun; stress on initial syllable) and “proTEST” (verb; stress on final) have contrasting stress patterns but the same phonemes (with predictable changes to the quality of the initial vowel depending on stress).
@@
This paper offers several novel contributions. First, a new speech dataset was automatically gathered without human annotation, using a pipeline built around deep learning systems. Second, we present a state-of-the-art model for lexical stress classification in disyllabic words and show its ability to generalize to novel words. Lastly, we employ LRP to analyze the spectral features these classifiers use to achieve high accuracy.
@@ II Dataset Construction
Previous studies on lexical stress often depended on stress labels assigned by human annotators [shahin2016automatic, korzekwa2020detection, 1415269] or were assigned based on the canonical stress for the orthographic word taken from pronunciation dictionaries such as the CMU dictionary [LI201828]. However, manual annotation is costly and can introduce inter-annotator inconsistency, and the use of a single citation form can yield errors for orthographically ambiguous minimal pair words (e.g., protest). Prior work also failed to explicitly document measures to prevent lexical overlap between training and evaluation splits, leaving open the possibility that identical word types were seen during both stages [shahin2016automatic, korzekwa2020detection, 1415269, LI201828]. Reported accuracy may therefore overestimate true generalization. To avoid these concerns, we created a novel dataset, using a fully automatic pipeline, without utilizing human annotations [^1] The complete dataset, including all audio samples, is publicly available at https://huggingface.co/datasets/MLSpeech/lexical_stress_dataset..
@@ II.1 Selection of English Disyllabic Words
We prompted ChatGPT 4o [hurst2024gpt] to generate 30 pairs of disyllabic words that form stress minimal pairs: pairs of words with the same spelling but different meaning depending on the primary stress placement. Five of these items could not be incorporated into other parts of our pipeline (the part-of-speech tagger, discussed below, failed to correctly annotate these forms). The remaining 25 pairs are address, conduct, content, contrast, convert, decrease, digest, export, extract, import, increase, insult, object, perfect, present, progress, project, protest, rebel, record, refuse, reject, subject, suspect, and transfer; the initial and final syllable stress variants of these words will be referred to as “minimal pairs.” We also asked ChatGPT 4o to generate a larger set of 250 disyllabic words that do not have stress minimal pairs (e.g., initial stress WALLet, final stress exTEND). After excluding one error, we retained 249 words (124 initial stress, 125 final stress). We refer to this as the set of “no minimal pair words.”
@@ II.2 Extraction of Words from Existing Corpora
The processing pipeline is shown in Figure 1. First, we used a forced alignment algorithm to align the transcript with the audio. Specifically we used Montreal Forced Aligner (MFA) [mcauliffe2017montreal] to align LibriSpeech and Supreme Court. MFA works at time-resolution of 10 millisecond and shown by rousso2024tradition to outperform recent end-to-end ASR-based aligners such as WhisperX [bain2023whisperx] and MMS [pratap2024scaling]. For TED-LIUM, which already includes word-level timestamps, we used the provided alignments. Importantly, unlike MFA-aligned corpora (LibriSpeech and Supreme Court), where word boundaries are adjacent (e.g., “hello 0–0.3 second”, “my 0.3–0.4”), the TED-LIUM timestamps may be non-adjacent (e.g., “hello 0–0.3”, “my 0.35–0.4”), leaving short pauses between words.
Once the start time of each word is obtained, a 0.5-second window is opened. This window size was chosen to ensure that both syllables of every disyllabic word were fully included and to maintain a uniform input length compatible with our CNN architecture. We retained only those whose total duration was less than 0.5 second and zero-padded them as needed to reach exactly 0.5 second. Each minimal pair word was further segmented into syllables using a syllabification algorithm [^2] The syllabified dictionary and syllabifier can be accessed at http://sourceforge.net/p/p2tk/.. In addition, phoneme-level timestamps derived from the forced aligner were incorporated.
@@
Overall, we had 7,446 samples (124 types) and 3,263 samples (125 types) for initial and final stress, respectively, in the no minimal pairs set, and 5,476 samples and 1,715 samples for initial and final stress, respectively, in the minimal pairs set, consisting of a total of 25 pairs.
@@ II.3 Phonetic Properties of Extracted Samples
To provide an initial verification that our procedures yielded sets of disyllabic words that contrasted in stress location, two well-known correlates of lexical stress were examined: amplitude (vowels in stressed syllables are louder than those in unstressed syllables) and duration (vowels in stressed stressed syllables are longer than those in unstressed syllables) [fry1955duration, fry1958experiments]. For each sample, we calculated the ratio of the properties of the vowels of the initial versus final syllables. These relative measures allow us to control for potential differences in speech rate and amplitude across samples. Our samples reliably showed the expected differences. Samples with initial stress had high ratios of amplitude (mean 0.6, s.d. 0.18) and duration (mean 0.53, s.d. 0.15), reflecting relatively loud, long vowels in stressed initial versus unstressed final syllables. In contrast, samples with final stress showed lower mean ratios for amplitude (mean 0.29, s.d. 0.14) and duration (mean 0.35, s.d. 0.12), reflecting relatively quiet, short vowels in unstressed initial versus stressed final syllables. To confirm that this is a reliable difference between sample types, we used bootstrap resampling (with 1,000 replicates) to estimate the 95% confidence interval for differences in mean ratios between initial versus final stress samples. These differences were significant (amplitude mean difference: 0.305; 95% CI (0.302,0.309); p<p<.0001; duration mean difference: 0.178, 95% CI (0.175,0.18), $p<$ .0001). This provides independent confirmation that the samples in this data set have contrasting stress patterns.
@@ III.2 Data Splitting
The training set was constructed using 201 out of the 249 word types from the no minimal pair words. Data augmentation techniques were employed to enlarge this set of examples. For each sample, we added a low-pass filtered version of the sample (cutoff frequency at 3000 Hz) as well as three versions of the sample mixed with noise. A sample of multi-talker babble from the VOiCES Corpus [richey2018voices] was combined with the original sample at three different levels (20 dB, 10 dB, and 3 dB SNR). These augmentations aimed to simulate real-world conditions, enhancing the robustness of our models. After the augmentation, we had 29,830 and 13,510 samples from the initial and final stress, respectively, corresponding to the 201 no minimal pair word types. The augmented dataset will be referred to as the “no minimal pairs train.”
The remaining 48 no minimal pair word types were divided into two lists of 24 words each, with one list used for the validation (i.e., hyper-parameter tuning; “no minimal pairs validation”) and the other for the test set (“no minimal pairs test”). The 25 minimal pairs words will be referred to as “minimal pairs test”. Note that two test sets were used (“no minimal pairs test” and “minimal pairs test”) as the minimal pairs set may contain errors; stress was determined solely by part-of-speech tagging, which may be inaccurate. In contrast, the tagging for the no minimal pair words is more reliable; as the word form is unambiguous, the lexically-determined stress pattern can be accurately identified. See Table 1 for dataset composition.
@@
minimal pairs test 25 5476 1715
@@ III.3 Training Process
The models were trained using standard procedures for CNNs. To stabilize training, we employed an Adam Optimizer [kingma2014adam] with a learning rate $5\times 10^{-4}$ and batch normalization. To address class imbalance (i.e., more initial versus final stress examples), we employed the focal loss function [lin2017focal], modifying the standard cross-entropy loss to focus more on hard-to-classify examples. Focal loss introduces a modulating factor $(1-p_{t})^{\gamma}$ to the cross-entropy objective, reducing the relative loss contribution of well-classified examples and amplifying that of misclassified ones. Formally, it is defined as:
@@
Each model was trained for up to 50 epochs, with early stopping triggered after 15 epochs without validation improvement. All experiments were conducted on 8 NVIDIA A40 GPUs.
@@ IV.1 Layer-wise Relevance Propagation
In our study, we apply this technique [kokhlikyan2020captum] to our lexical stress classifiers, backpropagating prediction to time-frequency bins in the input [^3] We used the implemented in the Captum library https://captum.ai/api/lrp.html.. This method identifies the regions of the spectrogram that have the most significant influence on the model’s classification of lexical stress. We specifically examine our best-performing VGG-based classifier, VGG16, and ResNet18 (for model performance comparisons see Section V and Table 2).
Figure 3: Further illustrative examples of spectrograms and corresponding attributions for the words “subject,” “address” and “refuse” with different primary stress location, using the VGG16 model. Vertical lines and annotation follow Figure 2. Panel (a)-(d) show spectrograms and corresponding $\mathbf{LRP_{CMP}}$ heatmaps for the word ”subject” with initial and final stress, respectively. Similarly, panels (e)-(l) display spectrograms and heatmaps of the words “address” and “refuse,” with different stress types. Note that the examples were drawn from datasets in which word segments were zero-padded to reach the fixed 0.5 second window length, as described in Section II.2.
@@ V Model Performance
Table 2: Classification accuracy for each model architecture on each stress type and test set.
@@ VI.2 Bounding Box and IOU Analysis
Overall, these findings suggest that the classifiers predict the location of stress by attending to features of the stressed syllable, in particular to stressed vowels. However, it is important to note that under all methods, a considerable number of pixels lie outside of the stressed vowel – consistent with the distribution of cues to lexical stress throughout the entire syllable. Furthermore, many pixels lie outside of the stressed syllable – consistent with a view that sees lexical stress as a “relational” property of stressed as compared to unstressed syllables.
@@ VI.3 Feature-Specific Relevance Analysis
Figure 5: Classification accuracy for each minimal pair word obtained with the VGG16 model. Each bar represents the proportion of correctly classified tokens for a given minimal pair word. Bars corresponding to words with vowel reduction are shown in dark blue, and those without vowel reduction are shown in light blue with hatching. The number of tokens for each word is indicated in parentheses. Words are ordered by accuracy.
@@ VII Discussion
To explore what allows CNNs to outperform more traditional approaches, we built several novel classifiers for lexical stress detection in disyllabic words, using LRP-based analyses to explore the spatiotemporal properties that drive model performance. A fully automatic pipeline was used to extract disyllabic words from read and spontaneous speech. CNNs trained on words without minimal pairs exhibited high performance on held-out no minimal pair words and words with minimal pairs. LRP analyses on the spectrogram input space for the CNNs revealed that the models were sensitive to acoustic properties in both stressed and unstressed syllables, but showed greater reliance on stressed syllables (particularly the stressed vowel). Analysis of our highest-performing model showed that $F_{1}$ of the stressed vowel exerted a stronger influence on predictions than other formants, but $F_{2}$ and, to a lesser extent, $F_{3}$ also contributed. Furthermore, these features failed to capture a substantial amount of variance in the LRP heatmap, suggesting additional features of the stressed vowel contribute to its predictions (e.g., spectral balance and pitch based on spacing of harmonics).
@@
The classifiers’ greater reliance on the properties of stressed versus unstressed syllables and more specifically stressed vowels versus other syllable regions is similar to the focus of the phonetics literature on stressed syllables and stressed vowels [cutler, gay1978physiological]. Critically, differing from some early phonetics studies, the classifier does not focus exclusively on these regions, but also attends to information in both syllables, including non-vowel portions [^4] It is possible that distributed nature of stress cues reflects errors in our forced aligner. However, given the substantial temporal range of LRP pixels (see, e.g., Figure 4(c)), it’s likely that the results are robust to small changes to segment boundaries.. The use of information in the unstressed syllable resonates with the relatively widespread use of relativized or normalized cues to lexical stress [cutler, vanHeuvenStress]. The use of information outside of the vowel is also consistent with phonetics research that examines duration and intensity of the whole syllable, the rime, or consonants [cutler, gordon2017acoustic]. More broadly, the use of consonantal information is consistent with the view that stress is not a property of vowels but of whole syllables, organized within more complex prosodic structures such as feet; these structures condition not only the properties of vowels but also surrounding consonants in stressed and unstressed syllables (see, e.g., [jensen2000against]). Future work should continue to develop this general line of analysis by examining variation in stress driven by larger sentential and discourse contexts (see vanHeuvenStress for a review and discussion).
@@
While stressed versus unstressed syllables are typically distinguished by full versus reduced vowel quality, some disyllabic English stress-pairs are pronounced with full vowels in both syllables. Figure 5 summarizes classification accuracy for each of the minimal pair words in the present study. Words whose unstressed syllable undergoes vowel reduction (e.g., PREsent vs. preSENT) are shown in dark blue, whereas words without vowel reduction (e.g., IMport vs. imPORT) are shown in light blue with hatching. Tagging of minimal pair words as “with” or “without” vowel reduction was based on a predefined list of English stress pairs [^5] The word list used to categorize minimal pairs as with and without vowel reduction is available at https://osf.io/5d4ks/overview. (n.b. the actual degree of reduction in particular productions may not precisely align with these assumptions). The figure reveals a clear performance gap: the models achieve substantially higher accuracy for words exhibiting vowel reduction than for those that do not. This pattern is consistent with the nature of the training data, which consisted primarily of no minimal pairs disyllabic words in which vowel reduction serves as an important acoustic correlate of stress [vanHeuvenStress]. It also supports the interpretation that formant information, particularly cues related to vowel reduction, plays a central role in model performance, while also confirming that the models are sensitive to additional cues in cases where vowel quality differences are minimal.
@@
More broadly, this work suggests that deep learning may offer an important complement to traditional approaches to phonetics. Rather than starting with highly controlled materials like minimal pairs elicited under laboratory conditions, our classifier was trained on “messy” data – uncontrolled materials, varying along many dimensions beyond the contrast of interest, drawn from speech produced by a number of talkers in a variety of conditions. When exposed to this (naturally occurring) variability, the classifiers were able to extract regularities that fit with the previous phonetics literature. With increasingly powerful interpretability tools, deep learning may help phoneticians better understand the acoustic properties associated with linguistic contrasts.

### v3

# How Does a Deep Neural Network Look at Lexical Stress in English Words?

@@ I Introduction
While deep learning based architectures have transformed many aspects of speech processing, those models often exhibit an nontransparent prediction process, obscuring how specific decisions are made. This work explores the application of one technique for shedding light on such architectures through the lens of a well-studied phenomenon in speech science and speech processing: lexical stress. Lexical stress refers to the prominence, accentuation, or emphasis of a particular syllable within a word. In many languages, lexical stress is crucial for communication as it can serve as the primary basis for a contrast in meaning. For example, in English, the words ”PROtest” (noun; stress on initial syllable) and ”proTEST” (verb; stress on final) have contrasting stress patterns but the same phonemes (with predictable changes to the quality of the initial vowel depending on stress).
@@
This paper offers several novel contributions. First, a new speech dataset was automatically gathered without human annotation, using a pipeline built around deep learning systems. Second, we present a state-of-the-art model for lexical stress classification in disyllabic words and show its ability to generalize to novel disyllabic words. Lastly, we employ LRP to analyze the spectral features these classifiers use to achieve high accuracy.
@@ II Dataset Construction
Previous studies on lexical stress often depended on stress labels assigned by human annotators [shahin2016automatic, korzekwa2020detection, 1415269] or were assigned based on the canonical stress for the orthographic word taken from pronunciation dictionaries such as the CMU dictionary [LI201828]. However, manual annotation is costly and can introduce inter-annotator inconsistency, and the use of a single citation form can yield errors for orthographically ambiguous minimal pair words (e.g., protest). Prior work also failed to explicitly document measures to prevent lexical overlap between training and evaluation splits, leaving open the possibility that identical word types were seen during both stages [shahin2016automatic, korzekwa2020detection, 1415269, LI201828]. Reported accuracy may therefore overestimate true generalization. To avoid these concerns, we created a novel dataset, using a fully automatic pipeline, without utilizing human annotations [mlspeech_lexical_stress_2025].
@@ II.1 Selection of English Disyllabic Words
We prompted ChatGPT 4o [hurst2024gpt] to generate 30 pairs of English disyllabic words that form stress minimal pairs: pairs of words with the same spelling but different meaning depending on the primary stress placement. Five of these items could not be incorporated into other parts of our pipeline (the part-of-speech tagger, discussed below, failed to correctly annotate these forms). The remaining 25 pairs are address, conduct, content, contrast, convert, decrease, digest, export, extract, import, increase, insult, object, perfect, present, progress, project, protest, rebel, record, refuse, reject, subject, suspect, and transfer; the initial and final syllable stress variants of these words will be referred to as ”minimal pairs.” We also asked ChatGPT 4o to generate a larger set of 250 English disyllabic words that do not have stress minimal pairs (e.g., initial stress WALLet, final stress exTEND). After excluding one error, we retained 249 words (124 initial stress, 125 final stress). We refer to this as the set of ”no minimal pair words.”
@@ II.2 Extraction of Words from Existing Corpora
The processing pipeline is shown in Figure 1. First, we used a forced alignment algorithm to align the transcript with the audio. Specifically we used Montreal Forced Aligner (MFA) [mcauliffe2017montreal] to align LibriSpeech and Supreme Court. MFA works at time-resolution of 10 millisecond and shown by rousso2024tradition to outperform recent end-to-end ASR-based aligners such as WhisperX [bain2023whisperx] and MMS [pratap2024scaling]. For TED-LIUM, which already includes word-level timestamps, we used the provided alignments. Importantly, unlike MFA-aligned corpora (LibriSpeech and Supreme Court), where word boundaries are adjacent (e.g., ”hello 0–0.3 second”, ”my 0.3–0.4”), the TED-LIUM timestamps may be non-adjacent (e.g., ”hello 0–0.3”, ”my 0.35–0.4”), leaving short pauses between words.
Once the start time of each word is obtained, a 0.5-second window is opened. This window size was chosen to ensure that both syllables of every disyllabic word were fully included and to maintain a uniform input length compatible with our CNN architecture. We retained only those whose total duration was less than 0.5 second and zero-padded them as needed to reach exactly 0.5 second. Each minimal pair word was further segmented into syllables using a syllabification algorithm [p2tk_syllabifier_2014]. In addition, phoneme-level timestamps derived from the forced aligner were incorporated.
@@
Overall, we had 7,446 samples (124 types) and 3,263 samples (125 types) for initial and final stress, respectively, in the no minimal pairs set, and 5,475 samples and 1,715 samples for initial and final stress, respectively, in the minimal pairs set, consisting of a total of 25 pairs.
@@ II.3 Phonetic Properties of Extracted Samples
To provide an initial verification that our procedures yielded sets of disyllabic words that contrasted in stress location, two well-known correlates of lexical stress were examined: amplitude (vowels in stressed syllables are louder than those in unstressed syllables) and duration (vowels in stressed syllables are longer than those in unstressed syllables) [fry1955duration, fry1958experiments]. For each sample, we calculated the ratio of the properties of the vowels of the initial versus final syllables. These relative measures allow us to control for potential differences in speech rate and amplitude across samples. Our samples reliably showed the expected differences. Samples with initial stress had high ratios of amplitude (mean 0.6, s.d. 0.18) and duration (mean 0.53, s.d. 0.15), reflecting relatively loud, long vowels in stressed initial versus unstressed final syllables. In contrast, samples with final stress showed lower mean ratios for amplitude (mean 0.29, s.d. 0.14) and duration (mean 0.35, s.d. 0.12), reflecting relatively quiet, short vowels in unstressed initial versus stressed final syllables. To confirm that this is a reliable difference between sample types, we used bootstrap resampling (with 1,000 replicates) to estimate the 95% confidence interval for differences in mean ratios between initial versus final stress samples. These differences were significant (amplitude mean difference: 0.305; 95% CI (0.302,0.309); p<p<.0001; duration mean difference: 0.178, 95% CI (0.175,0.18), $p<$ .0001). This provides independent confirmation that the samples in this data set have contrasting stress patterns.
@@ III.2 Data Splitting
The training set was constructed using 201 out of the 249 word types from the no minimal pair words. Data augmentation techniques were employed to enlarge this set of examples. For each sample, we added a low-pass filtered version of the sample (cutoff frequency at 3000 Hz) as well as three versions of the sample mixed with noise. A sample of multi-talker babble from the VOiCES Corpus [richey2018voices] was combined with the original sample at three different levels (20 dB, 10 dB, and 3 dB SNR). These augmentations aimed to simulate real-world conditions, enhancing the robustness of our models. After the augmentation, we had 29,830 and 13,510 samples from the initial and final stress, respectively, corresponding to the 201 no minimal pair word types. The augmented dataset will be referred to as the ”no minimal pairs train.”
The remaining 48 no minimal pair word types were divided into two lists of 24 words each, with one list used for the validation (i.e., hyper-parameter tuning; ”no minimal pairs validation”) and the other for the test set (”no minimal pairs test”). The 25 minimal pairs words will be referred to as ”minimal pairs test”. Note that two test sets were used (”no minimal pairs test” and ”minimal pairs test”) as the minimal pairs set may contain errors; stress was determined solely by part-of-speech tagging, which may be inaccurate. In contrast, the tagging for the no minimal pair words is more reliable; as the word form is unambiguous, the lexically-determined stress pattern can be accurately identified. See Table 1 for dataset composition.
@@
minimal pairs test 25 5475 1715
@@ III.3 Training Process
The models were trained using standard procedures for CNNs. To stabilize training, we employed an Adam Optimizer [kingma2014adam] with a learning rate $5\times 10^{-4}$ and batch normalization. Training was performed for up to 50 epochs, with early stopping triggered after 15 epochs without validation improvement. All experiments were conducted on 8 NVIDIA A40 GPUs.
As the training data contain substantially more initial than final stress examples, a potential concern is an initial stress bias, whereby a classifier could achieve relatively high accuracy by disproportionately predicting the majority class, rather than relying on acoustically relevant cues to lexical stress. Such a bias could obscure the extent to which the model genuinely learns stress-related acoustic patterns. To mitigate this possibility, we employed the focal loss function [lin2017focal], modifying the standard cross-entropy loss to focus more on hard-to-classify examples and to reduce the influence of the majority class during training. Focal loss (FL) introduces a modulating factor $(1-p_{t})^{\gamma}$ to the cross-entropy objective, reducing the relative loss contribution of well-classified examples and amplifying that of misclassified ones. Formally, it is defined as:
@@ IV.1 Layer-wise Relevance Propagation
In our study, we apply this technique to our lexical stress classifiers, backpropagating prediction to time-frequency bins in the input using the Captum interpretability library [kokhlikyan_captum_2020]. This method identifies the regions of the spectrogram that have the most significant influence on the model’s classification of lexical stress. We specifically examine our best-performing VGG-based classifier, VGG16, and ResNet18 (for model performance comparisons see Section V and Table 2).
Figure 3: Further illustrative examples of spectrograms and corresponding attributions for the words ”subject,” ”address” and ”refuse” with different primary stress location, using the VGG16 model. Vertical lines and annotation follow Figure 2. Panel (a)-(d) show spectrograms and corresponding $\mathbf{LRP_{CMP}}$ heatmaps for the word ”subject” with initial and final stress, respectively. Similarly, panels (e)-(l) display spectrograms and heatmaps of the words ”address” and ”refuse,” with different stress types. Note that the examples were drawn from datasets in which word segments were zero-padded to reach the fixed 0.5 second window length, as described in Section II.2.
@@ V Model Performance
Table 2: Classification accuracy for each model architecture on each stress type and test set (Chance-level performance is 50%).
@@ VI.2 Bounding Box and IOU Analysis
Overall, these findings suggest that the classifiers predict the location of stress by attending to features of the stressed syllable, in particular to stressed vowels. However, it is important to note that under all methods, a considerable number of pixels lie outside of the stressed vowel – consistent with the distribution of cues to lexical stress throughout the entire syllable. Furthermore, many pixels lie outside of the stressed syllable – consistent with a view that sees lexical stress as a ”relational” property of stressed as compared to unstressed syllables.
@@ VI.3 Feature-Specific Relevance Analysis
Figure 5: Classification accuracy of the VGG16 model on the minimal pairs test set, aggregated by vowel reduction in the unstressed syllable. Bars show mean per-word accuracy, averaged across minimal pairs with vowel reduction (left) and without vowel reduction (right). Error bars indicate $\pm$ standard error of the mean computed across words within each group. The number of test samples in each group is indicated by n. Dashed red horizontal line denotes chance-level (random baseline) accuracy at 50%.
@@ VII Discussion
To explore what allows CNNs to outperform more traditional approaches, we built several novel classifiers for lexical stress detection in English disyllabic words, using LRP-based analyses to explore the spatiotemporal properties that drive model performance. A fully automatic pipeline was used to extract disyllabic words from read and spontaneous speech. CNNs trained on words without minimal pairs exhibited high performance on held-out no minimal pair words and words with minimal pairs. LRP analyses on the spectrogram input space for the CNNs revealed that the models were sensitive to acoustic properties in both stressed and unstressed syllables, but showed greater reliance on stressed syllables (particularly the stressed vowel). Analysis of our highest-performing model showed that $F_{1}$ of the stressed vowel exerted a stronger influence on predictions than other formants, but $F_{2}$ and, to a lesser extent, $F_{3}$ also contributed. Furthermore, these features failed to capture a substantial amount of variance in the LRP heatmap, suggesting additional features of the stressed vowel contribute to its predictions (e.g., spectral balance and pitch based on spacing of harmonics).
@@
The classifiers’ greater reliance on the properties of stressed versus unstressed syllables and more specifically stressed vowels versus other syllable regions is similar to the focus of the phonetics literature on stressed syllables and stressed vowels [cutler, gay1978physiological]. Critically, differing from some early phonetics studies, the classifier does not focus exclusively on these regions, but also attends to information in both syllables, including non-vowel portions  [^1] It is possible that the distributed nature of stress cues reflects errors in our forced aligner. However, given the substantial temporal range of LRP pixels (see, e.g., Figure 4(c)), it’s likely that the results are robust to small changes to segment boundaries.. The use of information in the unstressed syllable resonates with the relatively widespread use of relativized or normalized cues to lexical stress [cutler, vanHeuvenStress]. The use of information outside of the vowel is also consistent with phonetics research that examines duration and intensity of the whole syllable, the rime, or consonants [cutler, gordon2017acoustic]. More broadly, the use of consonantal information is consistent with the view that stress is not a property of vowels but of whole syllables, organized within more complex prosodic structures such as feet; these structures condition not only the properties of vowels but also surrounding consonants in stressed and unstressed syllables (see, e.g., [jensen2000against]). Future work should continue to develop this general line of analysis by examining variation in stress driven by larger sentential and discourse contexts (see vanHeuvenStress for a review and discussion).
@@
While, in English, stressed versus unstressed syllables are typically distinguished by full versus reduced vowel quality, some disyllabic stress-pairs are pronounced with full vowels in both syllables. Figure 5 summarizes classification accuracy on the minimal pairs test set after aggregating all minimal-pair words into two groups based on the presence or absence of vowel reduction in the unstressed syllable. Because the task is binary (initial vs. final stress), chance-level performance corresponds to 50% accuracy. Tagging of minimal pair words as ”with” or ”without” vowel reduction was based on a predefined list of English stress pairs [bosker_lexical_stress_minimal_pairs_2025]. Words treated as lacking vowel reduction include import, insult, digest, increase, and transfer. This classification was not based on token-level acoustic annotation; consequently, the actual degree of reduction in particular productions may not precisely align with these assumptions. The figure reveals a clear performance gap: the models achieve substantially higher accuracy for words exhibiting vowel reduction than for those that do not. When vowel reduction is absent, mean accuracy falls to chance level. This pattern is consistent with the nature of the training data, which consisted primarily of no minimal pairs disyllabic words in which vowel reduction serves as an important acoustic correlate of stress [vanHeuvenStress]. It also supports the interpretation that formant information, particularly cues related to vowel reduction, plays a central role in model performance, while also confirming that the models are sensitive to additional cues in cases where vowel quality differences are minimal.
@@
More broadly, this work suggests that deep learning may offer an important complement to traditional approaches to phonetics. Rather than starting with highly controlled materials like minimal pairs elicited under laboratory conditions, our classifier was trained on ”messy” data – uncontrolled materials, varying along many dimensions beyond the contrast of interest, drawn from speech produced by a number of talkers in a variety of conditions. When exposed to this (naturally occurring) variability, the classifiers were able to extract regularities that fit with the previous phonetics literature. With increasingly powerful interpretability tools, deep learning may help phoneticians better understand the acoustic properties associated with linguistic contrasts.